# One Tool Is Enough: Reinforcement Learning for Repository-Level LLM Agents

### v4

@@ 4.4 Scaling Law of Tool-Calling
To assess the significance of tool-calling in RepoNavigator, we varied the maximum number of tool-calling turns and reported the results in Fig. 4. As shown in the figure, allowing more tool-calling turns consistently leads to improved performance for RepoNavigator, both before and after reinforcement learning (RL) training. In other words, these results empirically validate the scaling law of tool-calling in this context.
@@ 4.5 Influence on Issue Resolution
To evaluate the impact of different localization results on the final issue resolution performance, we test RepoNavigator against baselines on SWE-bench_Verified. We directly apply the repairing phrase of Agentless while replacing its localization front-end with other methods. Table.3 illustrates the results. Compared with baselines, RepoNavigator has the highest performance on issue resolution, while reinforcement learning improves its performance furthermore.
Table 3: We use Qwen2.5-14B-Instruct as the localization model, and use Qwen2.5-32B-Instruct as the repair model on SWE-bench_Verified.
@@
Agentless 5.28 10.12
LocAgent 2.65 13.01
RepoNavigator 12.00 14.74
RepoNavigator+RL 14.58 15.03
@@ 5.3 Impact on the Prediction Space
The access scope of a tool is defined as the complete set of files, symbols, and other resources that the tool can access within a repository. For a jump tool that navigates to symbol definitions, its access scope can be obtained by starting from a given entry point and recursively resolving all referenced symbols until no new definitions can be reached. Apparently, its access scope is significantly smaller than the full repository scope. Consequently, when computing the Intersection over Union (IoU) between the prediction set and the groundtruth set, using the jump tool results in a higher IoU, as depicted in Fig. 5. On the other hand, applying multiple repo-level retrivel tools results in the access scope equal to the whole repository scope.
When we start from the entry point and repeatedly apply jump—which retrieves the definition of each referenced symbol—we effectively traverse all symbols that are semantically activated by that entry point. Because every location that contributes to the issue must lie on some dependency path originating from the entry point, it is necessarily reachable through this recursive symbol-reference expansion. Therefore, the final access scope produced by exhaustive jump traversal is guaranteed to contain all locations that must be modified to resolve the issue.
@@ 5.4 Verification
To further verify this proposal, we change the tool set of RepoNavigator and conduct RL training with only the outcome reward. We add excessive tools which were frequently used in previous works (Chen et al., 2025; Ma et al., 2025; Jiang et al., 2025) and present the result in Table. 4. GetClass/GetFunc takes a class/function name as input and outputs the class/function definition. GetStruc takes no input and outputs the repository’s structure. The results clearly implies that additional tools do not increase model’s performance. This inspires researchers to develop less but more capable tools.
@@ 6 Conclusion
Extensive experiments across SWE-bench-Verified and SWE-bench-Pro demonstrate that RepoNavigator achieves state-of-the-art localization performance. We theoretically analyze the results, confirming that a single powerful tool, jointly optimized with reinforcement learning, can provide stronger robustness and more reliable multi-step reasoning than previous frameworks relying on multiple narrowly scoped tools.
Our findings highlight the importance of aligning agent tooling with real execution structure, and show that efficient reasoning-tool co-training can unlock substantial gains even for medium-sized open-source models. Future work will explore extending RepoNavigator from Python to more programming languages.
@@ Appendix A Detailed Illustration of Baselines
RepoSearcher (Ma et al., 2025) is an agent that first conducts file-level localization and then function-level localization, which aligns with CoSIL. RepoSearcher introduced the first training framework ToolTrain for localization agents, which is composed of distilling from a close-source model (Claude3.7-Sonnet in RepoSeacher) as warmup and reinforcement learning to further enhance the performance.
@@
Compared with all baselines, we are the first fully-automatic LLM agent, with no fixed workflow and no planetary prompt, and we are the first method trained directly from pretrained open-source LLMs without a close-source teacher model. Lastly, we only integrate a single yet powerful tool to the agent, which reduces compounding error and narrows the access scope of the agent.
@@ Appendix B Experimental Details
Jump GetClass GetFunc GetStruc Function-level File-level
Recall Precision Sample-F1 IoU Recall Precision Sample-F1 IoU
14.28 15.44 14.40 13.71 35.78 36.76 35.59 34.55
22.60 25.02 22.80 21.44 48.49 50.13 48.52 47.17
24.64 27.48 25.05 24.00 53.48 55.76 53.68 52.69
25.11 29.16 25.75 24.28 55.81 58.71 56.32 54.89
@@
When the response exceeds the maximum length, we clip and force the agent to stop, and we give zero as its score. When the agent exceeds the maximum tool-calling times (which is 12), we add ”You must not call tools anymore, and you must give the final answer” to the tool’s response. Most of the time, the agent will stop calling tools and generate the final response. If not, we force it to stop and give zero as its score. Note that when the maximum tool-calling times is not achieved and the final answer is generated, the agent loop will stop automatically. The aforementioned process is an automatic agentic framework, which allows the agent to explore in the environments with little constraints.
@@
It is a widespread concern that data leakage at the pre-training phrase threatens the validity of post-training methods. Nevertheless, we exclude this concern by results in Tabel. 2. The SWE-bench_Pro dataset was published in 2025, while the Qwen2.5 series were published in 2024. Moreover, we exclude the samples in the training dataset if the repository also appears in SWE-bench_Verified or SWE-bench_Pro.
@@
Appendix C Illustration of Training
We present the detailed pseudo code of the RL training process of RepoNavigator bellow. For simplicity, we include the entry point (normally a testing function) in issue qq.
Algorithm 1 RL process of RepoNavigator
@@ Appendix C Illustration of Training
Appendix D Threats to Validity
@@
Another limitation is that we only evaluate Python repositories in our experiments. This is because each language (C/C++, Java, etc.) has its unique language server, and we only succeed in implementing the language server of python. We will implement more language servers and validate our approach on more programing languages in the future.
Appendix E Case Study
In this section, we present the full trajectory of RepoNavigator on astropy_astropy-12907 from SWE-bench_Verified in Fig. E. We apply the default tool-calling prompt template of verl (Shen, 2024) and present an example. The system prompt and user prompt are presented in Fig. E. Noted, we do not present any process restrictions in our prompt, encouraging RepoNavigator to plan, call tools, and make decisions full-automatically. This is distinct with Agentless (which has a fixed workflow), LocAgent (which predefines a specific step-by-step workflow in its system prompt), CoSIL and RepoSearcher (which is half-automatic because some forced steps are added to the workflow besides the automatic multi-turns tool-calling conversations).
Figure E: The following pages are prompt and case study.
@@ Appendix E Case Study
{"type": "function", "function": {"name": "check", "description": "In the
@@
The ’file_path’ is the relevant path of where the symbol is called,
@@
For instance, if ’classA.functionB’ is what you want to check (which is called
in fileA.py), you should directly check ’functionB’ in ’fileA.py’.
@@
{"name": "check", "arguments": {"symbol": "separability_matrix", "file_path":
@@
The definition of symbol ’separability_matrix’ is:
@@
The ‘separability_matrix‘ function is defined in ‘astropy/modeling/separable.py‘
@@
To address this, we need to check the ‘_separable‘ function, which is likely
where the separability logic is implemented. Let’s check the ‘_separable‘
@@
{"name": "check", "arguments": {"symbol": "_separable", "file_path":
@@
The definition of symbol ’_separable’ is:
@@
        transform : ‘astropy.modeling.Model‘
@@
        return _coord_matrix(transform, ’left’, transform.n_outputs)
@@
The ‘_separable‘ function is defined in ‘astropy/modeling/separable.py‘. This
@@
how ‘_operators‘ dictionary handles the combination of separability matrices.
Let’s check the ‘_operators‘ dictionary.
@@
{"name": "check", "arguments": {"symbol": "_operators", "file_path":
@@
The definition of symbol ’_operators’ is:
_operators = {’&’: _cstack, ’|’: _cdot, ’+’: _arith_oper, ’-’: _arith_oper,
’*’: _arith_oper, ’/’: _arith_oper, ’**’: _arith_oper}
@@
The ‘_operators‘ dictionary is defined in ‘astropy/modeling/separable.py‘.
@@
the problem might lie in how the ‘_cstack‘ function handles the combination
@@
To address this, we need to check the ‘_cstack‘ function.
Let’s check the ‘_cstack‘ function.
@@
{"name": "check", "arguments": {"symbol": "_cstack", "file_path":
@@
The definition of symbol ’_cstack’ is:
@@
    cleft = _coord_matrix(left, ’left’, noutp)
@@
    cright = _coord_matrix(right, ’right’, noutp)
@@
The ‘_cstack‘ function is defined in ‘astropy/modeling/separable.py‘. This
@@
for the ’&’ operator. Given the issue with nested CompoundModels, it seems
@@
Specifically, we need to check how the ‘_coord_matrix‘ function handles
@@
Let’s check the ‘_coord_matrix‘ function.
@@
{"name": "check", "arguments": {"symbol": "_coord_matrix", "file_path":
@@
The definition of symbol ’_coord_matrix’ is:
@@
The ‘_coord_matrix‘ function is defined in ‘astropy/modeling/separable.py‘. This
@@
with the ‘_cstack‘ function, which is responsible for handling the combination
of separability matrices for the ’&’ operator. The function does not correctly

### v5

@@ 4.4 Scaling Law of Tool-Calling
To assess the significance of tool-calling in RepoNavigator, we varied the maximum number of tool-calling turns and reported the results in Fig. 4. As shown in the figure, allowing more tool-calling turns consistently leads to improved performance for RepoNavigator, both before and after reinforcement learning (RL) training. These results empirically validate the scaling law of tool-calling in this context.
@@ 4.5 Influence on Issue Resolution
To evaluate the impact of different localization results on the final issue resolution performance, we test RepoNavigator against baselines on SWE-bench_Verified. We directly apply the repairing phase of Agentless while replacing its localization front-end with other methods. Table.3 illustrates the results. Compared with baselines, RepoNavigator has the highest performance on issue resolution, while reinforcement learning improves its performance furthermore.
Table 3: Downstream influence: We use Qwen2.5-14B-Instruct on SWE-bench_Verified to evaluate issue resolution.
@@
Agentless 12.28 10.12
LocAgent 12.32 13.01
RepoSearcher 10.60 12.47
RepoNavigator 25.58 14.74
RepoNavigator+RL 29.23 15.03
@@ 5.3 Impact on the Prediction Space
The access scope of a tool is defined as the complete set of files, symbols, and other resources that the tool can access within a repository. For a jump tool that navigates to symbol definitions, its access scope can be obtained by starting from a given entry point and recursively resolving all referenced symbols until no new definitions can be reached. Apparently, its access scope is significantly smaller than the full repository scope. Consequently, when computing the Intersection over Union (IoU) between the prediction set and the groundtruth set, using the jump tool results in a higher IoU, as depicted in Fig. 5. On the other hand, applying multiple repo-level retrieval tools results in the access scope equal to the whole repository scope.
Starting from a designated entry point, the recursive application of the jump operation—which resolves the definitions of all referenced symbols—effectively constructs a transitive closure of all semantically activated symbols. Since any code location contributing to a specific issue must inherently reside on a dependency path originating from the entry point, it is logically reachable through this recursive expansion of symbol references. Consequently, the final access scope generated by an exhaustive jump traversal is guaranteed to encompass all locations requiring modification to resolve the issue, excluding only those introduced by dynamic, run-time changes to the repository (e.g. monkey-patching and string-based imports).
@@ 5.4 Verification
To further verify this proposal, we change the tool set of RepoNavigator and conduct RL training with only the outcome reward. We add excessive tools which were frequently used in previous works (Chen et al., 2025; Ma et al., 2025; Jiang et al., 2025) and present the result in Table. 4. Due to page limit, we present the full result in Table. 6. GetClass/GetFunc takes a class/function name as input and outputs the class/function definition. GetStruc takes no input and outputs the repository’s structure in the file system. The results clearly implies that additional tools do not increase model’s performance. When one tool is capable enough, more tools does not mean more benefits. This inspires researchers to develop less but more capable tools.
@@ 6 Conclusion
Extensive experiments across SWE-bench_Verified and SWE-bench_Pro demonstrate that RepoNavigator achieves state-of-the-art localization performance. We theoretically analyze the results, confirming that a single powerful tool, jointly optimized with reinforcement learning, can provide stronger robustness and more reliable multi-step reasoning than previous frameworks relying on multiple narrowly scoped tools. Future work will explore extending RepoNavigator from Python to more programming languages.
@@ Appendix A Detailed Illustration of Baselines
The comprehensive toolsets for all aforementioned methods are detailed in Table. 5. Furthermore, this section provides an in-depth discussion regarding the degree of autonomy exhibited by these approaches.
@@
RepoSearcher (Ma et al., 2025) is an agent that first conducts file-level localization and then function-level localization, which aligns with CoSIL. RepoSearcher introduced the first training framework ToolTrain for localization agents, which is composed of distilling from a closed-source model (Claude3.7-Sonnet in Reposearcher) as warmup and reinforcement learning to further enhance the performance.
@@
Compared with all baselines, we are the first fully-automatic LLM agent, with no fixed workflow and no planetary prompt, and we are the first method trained directly from pretrained open-source LLMs without a closed-source teacher model. Lastly, we only integrate a single yet powerful tool to the agent, which reduces compounding error and narrows the access scope of the agent.
@@ Appendix B Experimental Details
Benchmarks
The SWE-bench (Jimenez et al., 2023) benchmark series is designed to evaluate the capability of Large Language Models (LLMs) in resolving real-world GitHub issues. In the official evaluation protocol, models are provided solely with the problem statement and the historical repository state to generate a patch. However, in practical software engineering workflows, developers typically have access to at least one entry point for any given issue. To better simulate these real-world scenarios and enhance evaluation fidelity, we deviate from the official SWE-bench standard by providing an explicit entry point for each sample. Specifically, for both our proposed method and all baseline models, we supply the precise file path and the corresponding entry function.
@@
Hyperparameters for Baselines
For baseline methods, we use vLLM (Kwon et al., 2023) to deploy the corresponding model locally, and we set max_respons_length to 32768, enforce-eager to True. We deploy all models on 8 NVIDIA Tesla-A100-80G GPUs. We follow the default settings of vLLM for other hyperparameters.
@@
Jump GetClass GetFunc GetStruc Function-level
Recall Precision Sample-F1 IoU
14.28 15.44 14.40 13.71
22.60 25.02 22.80 21.44
24.64 27.48 25.05 24.00
25.11 29.16 25.75 24.28
@@
When the response exceeds the maximum length, we clip and force the agent to stop, and we give zero as its score. When the agent exceeds the maximum tool-calling times (which is 12), we add “You must not call tools anymore, and you must give the final answer” to the tool’s response. Most of the time, the agent will stop calling tools and generate the final response. If not, we force it to stop and give zero as its score. Note that when the maximum tool-calling times is not achieved and the final answer is generated, the agent loop will stop automatically. The aforementioned process is an automatic agentic framework, which allows the agent to explore in the environments with little constraints.
@@
It is a widespread concern that data leakage at the pre-training phase threatens the validity of post-training methods. Nevertheless, we exclude this concern by results in Tabel. 2. The SWE-bench_Pro dataset was published in 2025, while the Qwen2.5 series were published in 2024. Moreover, we exclude the samples in the training dataset if the repository also appears in SWE-bench_Verified or SWE-bench_Pro.
@@
Appendix C More Results
Ablation Study
We investigate the impact of the hybrid reward mechanism in Fig. 3 and explore the influence of SFT as a “cold start” phase for the Reinforcement Learning (RL) process. As illustrated in Table 8, our findings suggest that while SFT enhances initial model performance, it ultimately constrains the long-term potential of RL. By forcing the model to adhere to pre-generated trajectories, SFT improves basic task proficiency at the expense of autonomous exploration. Consequently, the results in Table 8 demonstrate that training the model directly with reinforcement learning yields the most superior performance.
Cost Analysis
We conduct a comparative analysis of token usage across multiple methodologies, utilizing Qwen2.5-14B-Instruct as the base model on SWE-bench_Pro. Empirical results (see Table. 8) indicate that RepoNavigator represents a superior agentic design in terms of cost-efficiency. While baseline approaches such as CoSIL and RepoSearcher incur substantial overhead by indexing all repository files—leading to significantly higher average input tokens—our proposed method optimizes token consumption by avoiding such exhaustive operations.
Table 7: Ablation study on SFT and RL.
SFT RL Function-level
Recall Precision Sample-F1 IoU
15.89 17.46 16.19 15.46
24.92 27.44 25.38 24.01
25.93 30.08 26.97 25.75
26.69 30.34 27.49 26.43
Table 8: Cost Analysis.
Method Avg. Tokens
Input Output Total
Agentless 19842.96 2360.65 22203.16
CoSIL 7937.05 504.84 8441.09
LocAgent 45879.32 2979.10 22203.16
RepoSearcher 23845.10 2809.89 26654.99
RepoNavigator 1726.32 948.93 1917.07
Evolution of Tool-calling Proficiency
We evaluate the tool-calling performance during training in Fig. 6, where tool rewards consistently improve. A potential side effect of integrating success rates into the GRPO reward function is the emergence of ”conservative” agents that inflate success metrics by suppressing tool usage. However, the results in Fig. 6(b) and Fig. 6(c) mitigate this concern, showing that successful tool-calling turns increase while overall turns remain stable or rise. We observe a particularly pronounced increase in total turns for the 7B model (Fig. 6(b)). Unlike the 14B and 32B models, which leverage their superior reasoning to reach targets with fewer, more precise invocations, the 7B model undergoes a more intensive exploration phase due to its initially lower tool-mastery (Fig. 6(a)).
(a) Tool Reward
(b) Overall Turns
(c) Successful Turns
Figure 6: Training dynamics of tool-calling capabilities. (a) There is a steady improvement in tool rewards as training proceeds. (b) Notably, the 7B model shows a growing frequency of tool usage over time, distinguishing its behavioral evolution from that of the 14B and 32B scales. (c) A sustained increase in successful tool-calling turns underscores the effectiveness of the optimization process.
Appendix D More Retated Works
Aside from open-source baselines we reproduced in our research, there are also agentic designs for repo-level tasks. PAGENT (Xue et al., 2025) analyzes the failure reasons of existing agents, and applies CFG creation and exploration to infer the type of a patch. MarsCode (Liu et al., 2024b) is a training-free multi-agent framework including a manager, a searcher, a tester, a reproducer, and editors. Trae (Gao et al., 2025) is a training-free workflow composed of patch generation, regression test, patch pruning and selection. Among all agentic frameworks, Trae has the highest rank on SWE-bench leaderboard. Linxi (Yang et al., 2025c) is a training-free multi-agent framework while adding the repository as a knowledge database, and retrieval-based methods are applied to extract knowledge from this database. These workflows are sophisticated, and they successfully integrate repo-level tools including language server, terminal, and search engine. However, their complexity hinders their potential for training. On the other hand, we remove unnecessary tools and workflows, preserving only the jump tool (built upon language server) to build a fully-automatic agent. Our design enables efficient training with reinforcement learning.
Appendix E Illustration of Training
We present the detailed pseudo code of the RL training process of RepoNavigator in Algorithm. 1. For simplicity, we include the entry point (normally a testing function) in issue qq.
Algorithm 1 Reinforcement learning process of RepoNavigator using GRPO
@@ Appendix E Illustration of Training
[…]
22:  end for
23:end for
[…]
Appendix F Threats to Validity
@@
Another limitation is that we only evaluate Python repositories in our experiments. This is because each language (C/C++, Java, etc.) has its unique language server, and we only succeed in implementing the language server of python. We will implement more language servers and validate our approach on more programming languages in the future.
Appendix G Case Study
In this section, we present the full trajectory of RepoNavigator on astropy_astropy-12907 from SWE-bench_Verified in Fig. G. We apply the default tool-calling prompt template of verl (Shen, 2024) and present an example. The system prompt and user prompt are presented in Fig. G. Noted, we do not present any process restrictions in our prompt, encouraging RepoNavigator to plan, call tools, and make decisions full-automatically. This is distinct with Agentless (which has a fixed workflow), LocAgent (which predefines a specific step-by-step workflow in its system prompt), CoSIL and RepoSearcher (which is half-automatic because some forced steps are added to the workflow besides the automatic multi-turns tool-calling conversations).
Figure E: The following pages are prompt and case study on astropy_astropy_12907 of SWE-bench_Verified.
@@ Appendix G Case Study
{"type": "function", "function": {"name": "jump", "description": "In the
@@
The "file_path" is the relevant path of where the symbol is called,
@@
For instance, if "classA.functionB" is what you want to check (which is called
in fileA.py), you should directly check "functionB" in "fileA.py".
@@
{"name": "jump", "arguments": {"symbol": "separability_matrix", "file_path":
@@
The definition of symbol "separability_matrix" is:
@@
The "separability_matrix" function is defined in "astropy/modeling/separable.py"
@@
To address this, we need to check the "_separable" function, which is likely
where the separability logic is implemented. Let’s check the "_separable"
@@
{"name": "jump", "arguments": {"symbol": "_separable", "file_path":
@@
The definition of symbol "_separable" is:
@@
        transform : "astropy.modeling.Model"
@@
        return _coord_matrix(transform, "left", transform.n_outputs)
@@
The "_separable" function is defined in "astropy/modeling/separable.py". This
@@
how "_operators" dictionary handles the combination of separability matrices.
Let’s check the "_operators" dictionary.
@@
{"name": "jump", "arguments": {"symbol": "_operators", "file_path":
@@
The definition of symbol "_operators" is:
_operators = {"&": _cstack, "|": _cdot, "+": _arith_oper, "-": _arith_oper,
"*": _arith_oper, "/": _arith_oper, "**": _arith_oper}
@@
The "_operators" dictionary is defined in "astropy/modeling/separable.py".
@@
the problem might lie in how the "_cstack" function handles the combination
@@
To address this, we need to check the "_cstack" function.
Let’s check the "_cstack" function.
@@
{"name": "jump", "arguments": {"symbol": "_cstack", "file_path":
@@
The definition of symbol "_cstack" is:
@@
    cleft = _coord_matrix(left, "left", noutp)
@@
    cright = _coord_matrix(right, "right", noutp)
@@
The "_cstack" function is defined in "astropy/modeling/separable.py". This
@@
for the "&" operator. Given the issue with nested CompoundModels, it seems
@@
Specifically, we need to check how the "_coord_matrix" function handles
@@
Let’s check the "_coord_matrix" function.
@@
{"name": "jump", "arguments": {"symbol": "_coord_matrix", "file_path":
@@
The definition of symbol "_coord_matrix" is:
@@
The "_coord_matrix" function is defined in "astropy/modeling/separable.py". This
@@
with the "_cstack" function, which is responsible for handling the combination
of separability matrices for the "&" operator. The function does not correctly